An aqueous solution of HNO3 has an osmotic pressure of 0.80 atm at 27 deg celcius. The density of the solution is 1.03 g/ml. Determine its vapor pressure
1
Expert's answer
2020-09-30T07:50:54-0400

T = 27 °C = 300 K

Osmotic pressure is expressed by the formula:

Π = CRT

0.80 = C (mol/L) × 0.0821 ((L×atm)/(mol×K)) × 300 K

C = 0.032 mol/L

1 L of water contains 55.55 moles

n1 = 55.55 mol

n2 = 0.032 mol

Relative lowering of vapor pressure:

"\frac{P_0 – P_s}{P_0} = \frac{n_2}{n_2+n_1}"

At 27 °C pure water pressure is 1 atm

"\frac{1 – Ps}{1} = \frac{0.032}{55.55 + 0.032}"

1 – Ps = 0.000575

Ps = 0.999425 atm

Answer: 0.999 atm